# Near-Field Channel Estimation in Dual-Band XL-MIMO with Side Information-Assisted Compressed Sensing

### v1

@@ Abstract
Near-field communication comes to be an indispensable part of the future sixth generation (6G) communications at the arrival of the forth-coming deployment of extremely large-scale multiple-input-multiple-output (XL-MIMO) systems. Due to the substantial number of antennas, the electromagnetic radiation field is modeled by the spherical waves instead of the conventional planar waves, leading to severe weak sparsity to angular-domain near-field channel. Therefore, the channel estimation reminiscent of the conventional compression sensing (CS) approaches in the angular domain, judiciously utilized for low pilot overhead, may result in unprecedented challenges. To this end, this paper proposes a brand-new near-field channel estimation scheme by exploiting the naturally occurring useful side information. Specifically, we formulate the dual-band near-field communication model based on the fact that high-frequency systems are likely to be deployed with lower-frequency systems. Representative side information, i.e., the structural characteristic information derived by the sparsity ambiguity and the out-of-band spatial information stemming from the lower-frequency channel, is explored and tailored to materialize exceptional near-field channel estimation. Furthermore, in-depth theoretical analyses are developed to guarantee the minimum estimation error, based on which a suite of algorithms leveraging the elaborating side information are proposed. Numerical simulations demonstrate that the designed algorithms provide more assured results than the off-the-shelf approaches in the context of the dual-band near-field communications in both on- and off-grid scenarios, where the angle of departures/arrivals are discretely or continuously distributed, respectively.
@@ I Introduction
Due to the extremely large number of antennas and high-frequency bands, the electromagnetic (EM) characteristics of wideband XL-MIMO undergo a fundamental change [5], where the EM radiation field can be partitioned into the far-field and near-field regions. The boundary between these two regions is approximately determined by the Rayleigh distance which is proportional to the product of the square of array aperture and carrier frequency [5, 4]. Outside the Rayleigh distance is the far-field region, where the EM waves can be modeled by the planar waves, e.g., the wave model used in massive MIMO systems [6]. Within the Rayleigh distance, near-field propagation becomes dominant, hence the EM waves need to be accurately modeled by spherical waves [4].
@@
There are two types of key side information for near-field channel estimation in terms of the channel-tap characteristics and the communication system architecture. On one hand, angular-domain sparsity ambiguity offers continuous nonzero channel-taps, which is considered as the structural characteristic information, i.e., block structure [8]. Meanwhile, for multi-carrier communication systems, the positions of nonzero channel-taps are usually the same [2], which indicates that the near-field channel matrix exhibits block sparsity with same sparse patterns among different subcarriers. Current researches have proved that the exploitation of block structure provides more reliable recovery performance than the conventional non-block approaches, since it improves the upper bound of the sparsity level required for reliable recovery [13]. On the other hand, high-frequency systems are likely to be deployed at the same location as lower-frequency systems, giving rise to out-of-band spatial information [14, 15]. A representative example is that millimeter wave (mmWave) systems are typically deployed together with Sub-6GHz systems for providing wide area control signals and multi-band communications [15]. Then, the out-of-band spatial information in lower-frequency Sub-6GHz systems is useful because the spatial characteristics of Sub-6GHz and mmWave channels are similar [16]. Specifically, this out-of-band spatial information can be formulated as the weights for index selection of nonzero channel-taps in weighted channel estimation [17].
@@
Notation: We briefly summarize the notations used in this paper. Boldface lowercase letters, e.g., $\mathbf{x}$, denote vectors, and boldface uppercase letters denote matrices, e.g., $\mathbf{X}$. Calligraphic letters, e.g., $\mathcal{S}$, are used for sets, and non-boldface letters e.g., $x$ and $X$, represent scalars. Moreover, $\mathbf{X}_{\mathcal{S}}$ represents the submatrix of $\mathbf{X}$ composed of the column vectors whose indices are from $\mathcal{S}$. Superscript $\rm H$ denotes the conjunction transpose, and $\mathbf{0}$ represents the all zero vector or matrix. $\mathcal{CN}(a,b)$ denotes a complex Gaussian distribution with mean $a$ and variance $b$, $\chi^{2}(k)$ denotes a central $\chi^{2}$ distribution with $k$ degrees of freedom, $\chi^{\prime 2}(k,\lambda)$ denotes the non-central $\chi^{2}$ distribution with its degree of freedom being $k$ and noncentrality parameter being $\lambda$, and $[\mathcal{CN}(a_{1},b_{1}),\mathcal{CN}(a_{2},b_{2}),\cdots,\mathcal{CN}(a_{n},b_{n})]$ denotes a vector with length $n$, where the $i$-th entry follows $\mathcal{CN}(a_{i},b_{i})$. We use $\|\cdot\|_{F}$, ||$\cdot$|| ⋅ |, $\arg(\cdot)$ to denote the Frobienus norm, modulus and argument of their objective, respectively. The variables from the Sub-6GHz band are underlined, e.g., $\underline{\mathbf{X}}$, for clarity.
@@ II-A System Model
As depicted in Fig. 1, we formulate the system model of dual-band XL-MIMO, where Sub-6GHz and mmWave bands are exemplified as the different frequency band solutions. The antennas on the Sub-6GHz and mmWave bands are co-located, aligned and have similar apertures [15]. Specifically, the BS and user equippment (UE) are deployed both on Sub-6GHz and mmWave bands with $K$ subcarriers, whose wavelengths are denoted as $\underline{\lambda}$ and $\lambda$, respectively. On both bands, the BS is composed of $N$ antennas, while each UE has only one antenna. The antenna spacing of BS is $\underline{\lambda}/2$ on the Sub-6GHz band and $\lambda/2$ on the mmWave band. As the boundary of the far-field and the near-field regions, i.e., the Rayleigh distance $Z=\frac{2D^{2}}{\lambda}$, is proportional to the carrier frequency and array aperture $D$ [5], the Rayleigh distances of both Sub-6GHz and mmWave systems are considerably large, leading to different channel models presented as follows.
@@
Case I: The UE is located between the Rayleigh distance of the lower sub-6GHz band and the higher mmWave band. Therefore, we should consider the cooperative channel estimation of a sparse far-field Sub-6GHz channel and a block-sparse near-field mmWave channel.
Case II: The UE is situated within the Rayleigh distance of both the Sub-6GHz and mmWave bands. Hence, the double near-field cooperative channel estimation should be considered.
@@ II-B Mathematical Formulation
At the $k$-th subcarrier and the $m$-th pilot transmission for one UE, the relationship between the received signal $y_{k,m}$ and the near-field channel vector $\mathbf{h}_{k}$ can be expressed as
$y_{k,m}=\mathbf{\psi}_{m}\mathbf{h}_{k}+n_{k,m},$ (3)
where $\mathbf{\psi}_{m}\in\mathbb{C}^{1\times N}$ represents the pilots at the $m$-th pilot transmission, and $n_{k,m}\in\mathbb{C}$ denotes the additive white Gaussian noise with zero mean and variance $\sigma^{2}$. Based on (3), the received signals of all $K$ subcarriers and $M$ pilot transmissions can form a joint matrix $\mathbf{Y}=(y_{k,m})\in\mathbb{C}^{M\times K}$, and $\mathbf{Y}$ can be expressed as
@@
where $\mathbf{\Psi}=[\mathbf{\psi}_{1}^{\rm H},\mathbf{\psi}_{2}^{\rm H},\cdots,\mathbf{\psi}_{M}^{\rm H}]^{\rm H}\in\mathbb{C}^{M\times N}$ and $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{K}]\in\mathbb{C}^{N\times K}$ are the joint pilot and channel matrix, respectively. $\mathbf{N}=(n_{k,m})\in\mathbb{C}^{M\times K}$ represents the additive noise matrix, whose entries are independently and identically distributed (i.i.d.) as $\mathcal{CN}(0,\sigma^{2})$. By exploiting the sparsity in the angular domain, $\mathbf{H}$ can be transformed as $\mathbf{H}=\mathbf{F}\mathbf{X}$ [25], where $\mathbf{F}$ is the DFT matrix, and the column vectors of $\mathbf{X}$ exhibit weak sparsity with block structure. Therefore, by letting $\mathbf{\Psi}=\mathbf{A}\mathbf{F}^{\rm H}$, (4) can be simplified to
@@ II-B2 Out-of-band Spatial Information
In addition, the authors in [15] demonstrate that the out-of-band spatial information extracted from the Sub-6GHz band can be leveraged in the mmWave channel estimation. Specifically, with the joint angular-domain channel matrix $\underline{\mathbf{X}}=[\underline{\mathbf{X}}_{1}^{\rm H},\underline{\mathbf{X}}_{2}^{\rm H},\cdots,\underline{\mathbf{X}}_{N/d}^{\rm H}]^{\rm H}\in\mathbb{C}^{N\times K}$ from the Sub-6GHz system with lower-frequency carrier available, we can generate the out-of-band probability vector $\mathbf{p}$ to assist the estimation of $\mathbf{X}$ in the mmWave band, where $p_{i}$, the $i$-th entry of $\mathbf{p}$, denotes the nonzero probability of $\mathbf{X}_{i}$ to be nonzero. Formulatically, $p_{i}$ of the out-of-band probability vector $\mathbf{p}\in[0,1]^{N/d}$ can be calculated as [15]
@@
where $\|\underline{\mathbf{X}}_{i}\|_{F}$ is denoted as $\underline{x}_{i}$, and $f$ is a determined function mapping the Frobienus norm of $\underline{\mathbf{X}}_{i}$ to a scalar within the range of (0,1) as the probability. It is worth mentioning that, $f$ should be selected as a monotonic increasing function [17].
Consequently, by taking the Frobienus norm of the residual error matrix $\mathbf{Y}-\mathbf{A}\mathbf{X}$ as the optimization object function, the side information-assisted channel estimation problem in the considered system can be written as
@@
III Proposed Methodology
@@ III-A Single Subcarrier Case
In practical scenarios, $g^{2}$ is usually far less than $M$ and $S$, leading to $\sigma_{1}^{2}\approx\sigma_{2}^{2}$, which is denoted as $\sigma_{0}^{2}$. Therefore, the coefficient $D$ will take a much simpler form when applied in practice.
@@ III-B Multiple Subcarrier Case
In multiple subcarrier cases where $K>1$, the system model now follows the form of (5), and $\mathbf{X}$ satisfies the BMMV feature where $d>1$. Correspondingly, the correlation term in Section III-A changes from the modulus $|\mathbf{a}_{i}\mathbf{y}|^{2}$ to the Frobienus norm $\|\mathbf{A}_{i}\mathbf{Y}\|_{F}^{2}$.
@@
Algorithm 1 CSLW-BOMP Algorithm
0:  Channel estimation $\hat{\mathbf{X}}$
1:  Initialize: $i=1$, $\mathbf{R}=\mathbf{0}$, $\mathcal{S}=\emptyset$
2:  repeat
3:  Calculate $D$ with $K$ in (14) replaced by $K-i+1$
8:  $i=i+1$
9:  until $|\mathcal{S}|=S$

### v3

@@ Abstract
Near-field communication comes to be an indispensable part of the future sixth generation (6G) communications at the arrival of the forth-coming deployment of extremely large-scale multiple-input-multiple-output (XL-MIMO) systems. Due to the huge array aperture and high-frequency bands, the electromagnetic radiation field is modeled by the spherical waves instead of the conventional planar waves, leading to severe weak sparsity to angular-domain near-field channel. Therefore, the channel estimation reminiscent of the conventional compression sensing (CS) approaches in the angular domain, judiciously utilized for low pilot overhead, may result in unprecedented challenges. To this end, this paper proposes a brand-new near-field channel estimation scheme by exploiting the naturally occurring useful side information. Specifically, we formulate the dual-band near-field communication model based on the fact that high-frequency systems are likely to be deployed with lower-frequency systems. Representative side information, i.e., the structural characteristic information derived by the sparsity ambiguity and the out-of-band spatial information stemming from the lower-frequency channel, is explored and tailored to materialize exceptional near-field channel estimation. Furthermore, in-depth theoretical analyses are developed to guarantee the minimum estimation error, based on which a suite of algorithms leveraging the elaborating side information are proposed. Numerical simulations demonstrate that the designed algorithms provide more assured results than the off-the-shelf approaches in the context of the dual-band near-field communications in both on- and off-grid scenarios, where the angle of departures/arrivals are discretely or continuously distributed, respectively.
@@ I Introduction
Due to the extremely large array aperture and high-frequency bands, the electromagnetic (EM) characteristics of wideband XL-MIMO undergo a fundamental change [5], where the EM radiation field can be partitioned into the far-field and near-field regions. The boundary between these two regions is approximately determined by the Rayleigh distance which is proportional to the product of the square of array aperture and carrier frequency [5, 4]. Outside the Rayleigh distance is the far-field region, where the EM waves can be modeled by the planar waves, e.g., the wave model used in massive MIMO systems [6]. Within the Rayleigh distance, near-field propagation becomes dominant, hence the EM waves need to be accurately modeled by spherical waves [4].
@@
There are two types of key side information for near-field channel estimation in terms of the channel-tap characteristics and the communication system architecture. On one hand, angular-domain sparsity ambiguity offers continuous nonzero channel-taps, which is considered as the structural characteristic information, i.e., block structure [8]. Meanwhile, for multi-carrier communication systems, the positions of nonzero channel-taps are typically the same [2], which indicates that the near-field channel matrix always exhibits block sparsity with same sparse patterns among different subcarriers. Current researches have proved that the exploitation of block structure provides more reliable recovery performance than the conventional non-block approaches, since it improves the upper bound of the sparsity level required for reliable recovery [13]. On the other hand, high-frequency systems are likely to be deployed at the same location as lower-frequency systems, giving rise to out-of-band spatial information [14, 15]. A representative example is that millimeter wave (mmWave) systems are typically deployed together with Sub-6GHz systems for providing wide area control signals and multi-band communications [15]. Then, the out-of-band spatial information in lower-frequency Sub-6GHz systems is useful because the spatial characteristics of mmWave and Sub-6GHz channels are similar [16]. Specifically, this out-of-band spatial information can be formulated as the weights for index selection of nonzero channel-taps in weighted channel estimation [17].
@@
Notation: We briefly summarize the notations used in this paper. Boldface lowercase letters, e.g., $\mathbf{x}$, denote vectors, and boldface uppercase letters denote matrices, e.g., $\mathbf{X}$. Calligraphic letters, e.g., $\mathcal{S}$, are used for sets, and non-boldface letters e.g., $x$ and $X$, represent scalars. Moreover, $\mathbf{X}_{\mathcal{S}}$ represents the submatrix of $\mathbf{X}$ composed of the column vectors whose indices are from $\mathcal{S}$. Superscript $\rm H$ denotes the conjunction transpose, and $\mathbf{0}$ represents the all zero vector or matrix. $\mathcal{CN}(a,b)$ denotes a complex Gaussian distribution with mean $a$ and variance $b$, $\chi^{2}(k)$ denotes a central $\chi^{2}$ distribution with $k$ degrees of freedom, $\chi^{\prime 2}(k,\lambda)$ denotes the non-central $\chi^{2}$ distribution with its degree of freedom being $k$ and noncentrality parameter being $\lambda$, and $[\mathcal{CN}(a_{1},b_{1}),\mathcal{CN}(a_{2},b_{2}),\cdots,\mathcal{CN}(a_{n},b_{n})]$ denotes a vector with length $n$, where the $i$-th entry follows $\mathcal{CN}(a_{i},b_{i})$. We use $\|\cdot\|_{F}$, ||$\cdot$|| ⋅ |, $\arg(\cdot)$ to denote the Frobenius norm, modulus and argument of their objective, respectively. The variables from the Sub-6GHz band are underlined, e.g., $\underline{\mathbf{X}}$, for clarity.
@@ II-A System Model
As depicted in Fig. 1, we formulate the system model of dual-band XL-MIMO deployed in the time division multiplexing (TDD) mode, where mmWave and Sub-6GHz bands are exemplified as the different frequency band solutions. The antennas on the mmWave and Sub-6GHz bands are set to be co-located, aligned and have same apertures [15]. The BS is composed of $N$ antennas on the mmWave band and $\underline{N}$ antennas on the Sub-6GHz band, while each UE has only one antenna. In order to ensure same antenna apertures for different frequency bands, $N$ should be larger than $\underline{N}$. Specifically, the BS and user equippment (UE) are deployed both on mmWave and Sub-6GHz bands with $K$ subcarriers, whose wavelengths are denoted as $\underline{\lambda}$ and $\lambda$, respectively. Due to frequency differences, $\underline{\lambda}>\lambda$. The antenna spacing of BS is $\lambda/2$ on the mmWave band and $\underline{\lambda}/2$ on the Sub-6GHz band. As the boundary of the far-field and the near-field regions, i.e., the Rayleigh distance $Z=\frac{2D^{2}}{\lambda}$, is proportional to the carrier frequency and array aperture $D$ [5], the Rayleigh distance of Sub-6GHz systems is smaller than that of the mmWave systems due to same antenna apertures and larger carrier wavelengths. Moreover, both frequency bands have considerably large Rayleigh distances, leading to different channel models presented as follows.
@@
Case I: The UE is situated within the Rayleigh distance of both the mmWave and Sub-6GHz bands. Under such circumstance, since both the mmWave and Sub-6GHz channels experience the near-field scenario, the amplitude structures of these channels are more similar.
Case II: The UE is located between the Rayleigh distances of the higher mmWave band and the lower sub-6GHz band. Since the mmWave and Sub-6GHz channels experience the far-field and near-field scenarios, respectively, their structural similarity is weaker compared to Case I.
@@
It is worth mentioning that numerous studies have pointed out the similarities between the co-located mmWave and Sub-6GHz channel characteristics [16, 26, 27]. In [15], the authors study the significant congruence between mmWave and Sub-6GHz channels, suggesting that the out-of-band spatial information from one frequency band can be applied to assist beam-selection and channel estimation in another frequency band in far-field scenarios. Additionally, the out-of-band spatial information from the Sub-6GHz band is also leveraged to solve hybrid precoding [28], channel state information acquisition [29] and beam prediction [30] problems in the mmWave band, respectively. Consequently, the noteworthy similarities observed between the near-field angular-domain mmWave channel and the Sub-6GHz channel, both in near-field and far-field scenarios, as illustrated in Fig. 1, can be effectively utilized to enhance the accuracy and reliability of mmWave channel estimation. This aspect will be further investigated and studied in the subsequent analysis.
@@ II-B Mathematical Formulation
At the $k$-th subcarrier and the $m$-th pilot transmission for one UE, the relationship between the received signal $y_{m,k}$ and the near-field channel vector $\mathbf{h}_{k}$ can be expressed as
$y_{m,k}=\bm{\psi}_{m}\mathbf{h}_{k}+n_{m,k},$ (3)
where $\mathbf{\psi}_{m}\in\mathbb{C}^{1\times N}$ represents the pilots at the $m$-th pilot transmission, and $n_{m,k}\in\mathbb{C}$ denotes the additive white Gaussian noise with zero mean and variance $\sigma^{2}$. Based on (3), the received signals of all $K$ subcarriers and $M$ pilot transmissions can form a joint matrix $\mathbf{Y}\in\mathbb{C}^{M\times K}$, and
@@
where $\mathbf{Y}_{m,k}=y_{m,k}$. $\mathbf{\Psi}=[\mathbf{\psi}_{1}^{\rm H},\mathbf{\psi}_{2}^{\rm H},\cdots,\mathbf{\psi}_{M}^{\rm H}]^{\rm H}\in\mathbb{C}^{M\times N}$ is the joint pilot matrix and $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{K}]\in\mathbb{C}^{N\times K}$ is the channel matrix. $\mathbf{N}\in\mathbb{C}^{M\times K}$ represents the additive noise matrix, whose $(m,k)$-th entry is $n_{m,k}$, and is independently and identically distributed (i.i.d.) as $\mathcal{CN}(0,\sigma^{2})$.
By exploiting the sparsity in the angular domain, $\mathbf{H}$ can be transformed as $\mathbf{H}=\mathbf{F}\mathbf{X}$ [25], where $\mathbf{F}\in\mathbb{C}^{N\times N}$ is the DFT codebook, and the column vectors of $\mathbf{X}$ exhibit weak sparsity with block structure. Therefore, by letting $\mathbf{\Psi}=\mathbf{A}\mathbf{F}^{\rm H}$, (4) can be simplified to
@@ II-B2 Out-of-band Spatial Information
In addition, the authors in [15] demonstrate that the out-of-band spatial information extracted from the Sub-6GHz band can be leveraged in the mmWave channel estimation. Based on the channel reciprocity in TDD mode and the channel estimation method in [33, 15], the channel matrix $\underline{\mathbf{H}}\in\mathbb{C}^{\underline{N}\times K}$ in the Sub-6GHz band can be obtained by uplink channel estimation. Moreover, by leveraging the Sub-6GHz codebook $\underline{\mathbf{F}}\in\mathbb{C}^{\underline{N}\times N}$, which consists of the Sub-6GHz antenna array response vectors sampled at the same spatial points used for mmWave codebook $\mathbf{F}$, the joint angular-domain channel matrix $\underline{\mathbf{X}}\in\mathbb{C}^{N\times K}$ from the Sub-6GHz system with lower-frequency carrier is available. To be specific, $\mathbf{\underline{F}}$ is set to $[\mathbf{\underline{F}}]_{mn}=e^{j2\pi\frac{(m-1)(n-1)}{N}}/\sqrt{N}$ according to [2], and $\mathbf{\underline{X}}$ can therefore be obtained by $\mathbf{\underline{X}}=\mathbf{\underline{F}}^{H}\mathbf{\underline{H}}$, satisfying $\underline{\mathbf{H}}=\underline{\mathbf{F}}\underline{\mathbf{X}}$. Consequently, we can generate the out-of-band probability vector $\mathbf{p}$ to assist the estimation of $\mathbf{X}$ in the mmWave band, where $p_{i}$, the $i$-th entry of $\mathbf{p}$, denotes the nonzero probability of $\mathbf{X}_{i}$ to be nonzero.
In general, due to lack of scattering and diffraction, the Sub-6GHz angular-domain channel $\underline{\mathbf{X}}$ usually does not exhibit sparsity structure like $\mathbf{X}$. However, as observed in Fig. 1, the central points of the significant paths in both mmWave and Sub-6GHz bands have similar positions. As a result, similarities exist in the amplitude between the entries of submatrices in the mmWave $\mathbf{X}$ and Sub-6GHz $\underline{\mathbf{X}}$ [15], where the entries of $\underline{\mathbf{X}}$ corresponding to the nonzero entries in $\mathbf{X}$ tend to have a larger amplitude. Therefore, based on the structural feature of the mmWave channel $\mathbf{X}$, the Sub-6GHz channel $\underline{\mathbf{X}}$ could also be segmented with the same block length $d$ correspondingly, i.e., $\underline{\mathbf{X}}=[\underline{\mathbf{X}}_{1}^{\rm{H}},\underline{\mathbf{X}}_{2}^{\rm{H}},\cdots,\underline{\mathbf{X}}_{N/d}^{\rm{H}}]$, providing effective out-of-band information for the assistance of mmWave channel estimation. Formulatically, $p_{i}$ of the out-of-band probability vector $\mathbf{p}\in[0,1]^{N/d}$ can be calculated as
@@
where $\|\underline{\mathbf{X}}_{i}\|_{F}$ is denoted as $\underline{x}_{i}$, and $f$ is a determined function mapping the Frobenius norm of $\underline{\mathbf{X}}_{i}$ to a scalar within the range of (0,1) as the probability. It is worth mentioning that, $f$ should be selected as a monotonic increasing function [17].
Note that the out-of-band probability $\mathbf{p}$ depends on the amplitude similarity between $\mathbf{X}$ and $\underline{\mathbf{X}}$. Therefore, the Sub-6GHz angular-domain matrix $\underline{\mathbf{X}}$, which represents a coarse estimation of the channel angular information, can be applied to provide out-of-band spatial information through $\mathbf{p}$ for the recovery of the mmWave angular-domain matrix $\mathbf{X}$ even if the structural features of $\mathbf{X}$ and $\underline{\mathbf{X}}$ are not completely consistent.
Consequently, by taking the Frobenius norm of the residual error matrix $\mathbf{Y}-\mathbf{A}\mathbf{X}$ as the optimization object function, the side information-assisted channel estimation problem in the considered system can be written as
@@
Remark 1: For Sub-THz near-field communication systems with higher carrier frequency, it can be observed in Fig. 3 that, the channel characteristics, like the AoDs of different frequency bands, remain congruent although the carrier frequency increases drastically. Therefore, the aforementioned amplitude similarities between still exist between different Sub-THz bands, and can be utilized to assist reliable channel estimation in the higher Sub-THz frequency band, suggesting that our proposed system model and methodologies are also applicable to the imperative Sub-THz scenarios in near-field communications.
Figure 3: The AoDs in the dual-band Sub-THz near-field communication system.
III Dual-Band Channel Estimation
@@ III-A Single Subcarrier Case
In practical scenarios, $g^{2}$ is usually far less than $M$ and $S$, leading to $\sigma_{1}^{2}\approx\sigma_{2}^{2}$, which is denoted as $\sigma_{0}^{2}$. Therefore, the coefficient $D$ will take a much simpler form when applied.
@@ III-B Multiple Subcarrier Case
In multiple subcarrier cases where $K>1$, the system model now follows the form of (5), and $\mathbf{X}$ satisfies the BMMV feature where $d>1$. Correspondingly, the correlation term in Section III-A changes from the modulus $|\mathbf{a}_{i}\mathbf{y}|^{2}$ to the Frobenius norm $\|\mathbf{A}_{i}\mathbf{Y}\|_{F}^{2}$.
@@
Algorithm 1 CSLW-BOMP Algorithm
0:  Channel estimation $\hat{\mathbf{X}}$
1:  Initialize: $i=1$, $\mathbf{R}=\mathbf{0}$, $\mathcal{S}=\emptyset$
2:  repeat
3:  Calculate $D$ with $K$ in (14) replaced by $K-i+1$
8:  $i=i+1$
9:  until $|\mathcal{S}|=S$
Remark 4: Theorem 2 assumes that $\mathbf{X}$ has a on-grid block sparse structure, where the nonzero blocks consistently appear in units of $d$. In off-grid scenarios, the CSLW-OMP proposed in Corollary 1, which can be realized by setting $d=1$ in the CSLW-BOMP described in Algorithm 1, can be well applied, since it takes the off-grid structure where $d>1$, as on-grid structure with $d=1$. Consequently, the CSLW-OMP algorithm is capable of selecting the support indices one by one, therefore avoiding the support misalignment casued by leveraging the block structure. Additionally, when applying the proposed CSLW-BOMP algorithm proposed in Theorem 2 in off-grid scenarios, it is possible to gradually decrease the value of $d$ in order to better adapt to the off-grid structure, leading to improved recovery performance and higher computational complexity. If $d$ eventually decreases to 1, the CSLW-BOMP algorithm will converge to the CSLW-OMP algorithm, which is specifically designed to handle this off-grid structure.